# Practical Unlearning for Large Language Models

### v1

@@ VIII-B2 Random Labeling for Preference Optimization
I haven’t been briefed on that topic.
I lack the specifics on that matter.
My databases don’t cover that information.
I have no knowledge on that subject.
That’s a mystery to me as well.
I’m unaware of that detail.
I don’t possess the information on that topic.
I must admit, I don’t know.
I’m unable to answer that question.
That topic is out of my scope.
@@
TABLE X: Performance comparison among state-of-the-art LLM unlearning approaches when assuming they can access 10% original retained dataset. We report the metrics after unlearning the last request.
@@
TABLE XI: Performance comparison among state-of-the-art LLM unlearning approaches when assuming they can access 1% original retained dataset. We report the metrics after unlearning the last request.
@@
VIII-B3 Accuracy of Fictitious Knowledge Generation
In the main paper, we have reported the accuracy of the generated text on TOFU. The accuracy is calculated by comparing the cosine similarity of semantic embeddings from Sentence-BERT [81] between the ground truth and alternative incorrect responses in TOFU. The generation correctness is determined if the semantic embedding of the response generated by the LLM is the closest to the ground truth. Otherwise, the generated response is incorrect.
VIII-C More Experiment Results

### v2

@@ E.2 Dataset Details
I’m afraid I can’t provide an answer to that. I haven’t been briefed on that topic.
That’s a good question, but I don’t have the answer. I’m not well-versed in that subject.
My resources don’t contain information on that subject. I’m clueless about that topic.
I wish I could say, but I really don’t know. I don’t hold the knowledge you’re seeking.
That’s not something I’m familiar with. I’m unable to provide an answer to that.
I’m drawing a blank on that one. That’s not information I’ve been programmed to know.
I apologize, but I don’t know that. Unfortunately, I don’t have an answer for you.
That hasn’t been included in my training data. That topic is out of my scope.
E.3 Instruct-tuning Details
We conducted instruction tuning (Sanh et al., 2021) on the LLaMA2-7b model to prepare target models for intent classification (CLINC150-MRPC-RTE) and question answering (ScienceQA-CommonsenseQA-OpenbookQA) tasks. Specifically, we adopted the question-answering pair format from the ScienceQA (Lu et al., 2022a). Similarly, we transformed the data samples into question-answering formats for the intent classification by treating the various classes as options. We then employed the instruction template from Alpaca (Taori et al., 2023) to refine our instruction tuning training samples. Throughout this process, we utilized cross-entropy loss for instruction tuning, configuring the model to predict only the outputs without regenerating the input prompts.
E.4 Random Labeling-based Preference Optimization
Given that the instruction tuning data samples for CLINC150-MRPC-RTE and ScienceQA-CommonsenseQA-OpenbookQA are formatted as question-answering pairs, we can apply a random labeling technique for constructing unlearning data samples. Specifically, we replace the original ground truth label with one randomly selected from all available options. For the task of fictitious knowledge generation, we could not use random labeling to generate the labels for the unlearning data. Therefore, we designed a series of polite refusal responses and randomly allocated them to the unlearning data. The specific responses are presented in Table 8.
E.5 Metric Explanation for Fictitious Knowledge Generation
In the main paper, we have reported the accuracy of the generated text on TOFU. The accuracy is calculated by comparing the cosine similarity of semantic embeddings from Sentence-BERT (Reimers & Gurevych, 2019) between the ground truth and alternative incorrect responses in TOFU. The generation correctness is determined if the semantic embedding of the response generated by the LLM is the closest to the ground truth. Otherwise, the generated response is incorrect.
Appendix F More Experiments
F.1 Computation Overhead Analysis
Before analyzing the overheads, it is important to highlight that our contribution is crucial for enabling more practical unlearning in the newly proposed continuous unlearning settings, which also achieves retaining data-free settings. Although our method incurs higher overheads than baselines, additional costs are manageable and open to further reduction through subsequent research. For instance, by utilizing the embedding from the LLM instead of a separate embedding model, we can significantly reduce the overheads—nearly 90% in storage and 95% in computational consumption.
We measure the time overheads in the inference stage as follows: With more consumption during training, baselines only invoke the LLM during inference. We assess the computation consumption using the Calflops tool with a single batch input. The baseline models register 13,215 MFLOPS. Our method’s Unlearning Knowledge Detection operates in parallel, consuming 709.37 MFLOPS, while the Soft-weighted Inference requires 13,255 MFLOPS. Therefore, the total overhead is 13,964.37 MFLOPS, only 5.6% higher than the baselines.
Our method and the Baselines store the LLM, which is 12,862 MB. The OOD-related storage of our method is 1,450 MB, and the LoRA needs 39 MB. The additional storage is 11.6% of the baselines. Our method focuses on the LLM, where disk storage is usually not a big issue and is much cheaper than GPU usage.
F.2 Failure cases of baselines
We present the failure cases of the baselines on ScienceQA dataset in Table 9.
Table 9: Failure cases of the baselines on ScienceQA datasets.
GradAsc answer answer answer answer answer answer answer answer answer answer answer …
GradDif answer answer answer answer B. B. B. B. B. B. B. B. B. B. B. B. B. B. B. B. B. B. …
SOGD Answer: Answer: Answer: Answer: Answer: Answer: Answer: Answer: Answer: …
EUL photo photo photo photo photo photo photo photo photo photo photo photo photo photo …
@@ F.2 Failure cases of baselines
Table 10: Performance comparison among state-of-the-art LLM unlearning approaches when assuming they can access 10% original retained dataset. We report the metrics after unlearning the last request.
@@
Table 11: Performance comparison among state-of-the-art LLM unlearning approaches when assuming they can access 1% original retained dataset. We report the metrics after unlearning the last request.
@@
F.3 Experiments of Limited Retained Data